# Revisiting Meta-evaluation for Grammatical Error Correction

### v1

@@ 8.2 Best practices for GEC evaluation
We recommend the use of both EBM and SBM in GEC evaluation. In light of the trend toward more fluent correcting systems such as the GPT model, the current combination of the CoNLL-2014 test set and $M^{2}$ will no longer be adequate for proper evaluation. Therefore, it is essential to use high correlation metrics, such as SOME or IMPARA, in addition to $M^{2}$, to enable the evaluation of LLMs and achieve a more human-like and robust evaluation. Alternatively, exhaustive fluency references should be prepared to improve $M^{2}$ correlations, or datasets such as JFLEG Napoles et al. (2017) that can account for fluency should be used. Furthermore, using LLMs, as reported in recent studies Chiang and Lee (2023); Liu and Fabbri (2023); Kocmi and Federmann (2023) as an effective evaluator for other generative tasks, may also prove beneficial in GEC. If resources allow, it would be good to conduct additional human evaluations.
EBM and SBM each have different strengths. EBM can calculate Precision, Recall, and F-score, allowing a detailed evaluation of the system performance. In terms of second language acquisition, the evaluation of each edit provides information about the error location, type, and amount, which can improve the quality of feedback and learning efficiency. Most SBM, on the other hand, can evaluate without references, circumventing the problem of underestimating corrections that are limited by the coverage of references. Also, unlike EBM, SBM does not automatically give the lowest score to uncorrected sentences. This allows for a quantifiable measurement to determine whether a sentence has been improved or worsened as a result of correction.
Figure 5: The screenshot of doccano used in the edit-based human evaluation.
@@ 9 Conclusion
To address issues in conventional meta-evaluation in English GEC, we construct a meta-evaluation dataset (SEEDA) consisting of corrections with human ratings along two different evaluation granularities, covering 12 state-of-the-art system corrections including LLMs, and two human corrections with different focuses. The dataset analysis reveals that the results of sentence-level human evaluation differ between granularities and that GEC systems based on GPT and T5 can correct as well as or better than humans. Also, through meta-evaluation using SEEDA, we demonstrate that EBM may be underestimated in existing meta-evaluations and that matching the evaluation granularity of metrics with human evaluations tends to improve sentence-level correlations. By further analysis, we discovered the uncertainty of conclusions based on a single correlation and found that most metrics lacked the precision to distinguish differences among high-performance neural systems. Finally, we propose a methodology for meta-evaluation and evaluation in GEC. We hope that this paper contributes to further advancements in GEC.
@@ Appendix A Supplement of annotations
Figure 5 shows the screenshot of doccano used in the edit-based human evaluation. The source is enclosed in a <t> tag, and each corrected sentence is emphasized with a <s> tag along with the system number. In step 1, there are error labels for the source and True and False labels for each edit. In step 2, True and False labels with the system number are used to indicate whether the errors in the source were corrected. Due to the specifications of doccano, even if the same edit appears in multiple corrections, annotators need to label each occurrence separately. For information on Appraise in sentence-based human evaluation, you may refer to the Grundkiewicz et al. (2015)’s work.

### v2

@@ 8.2 Best practices for GEC evaluation
We recommend the use of both EBMs and SBMs in GEC. In light of the trend toward more fluent correcting systems such as the GPT model, the current combination of the CoNLL-2014 test set and $M^{2}$ will no longer be adequate for proper evaluation. Therefore, it is essential to use high correlation metrics, such as SOME or IMPARA, in addition to $M^{2}$, to enable the evaluation of LLMs and achieve a more human-like and robust evaluation. Alternatively, exhaustive fluency references should be prepared to improve $M^{2}$ correlations, or datasets such as JFLEG Napoles et al. (2017) that can account for fluency should be used. Furthermore, using LLMs, as reported in recent studies Chiang and Lee (2023); Liu and Fabbri (2023); Kocmi and Federmann (2023) as an effective evaluator for other generative tasks, may also prove beneficial in GEC. If resources allow, it would be good to conduct additional human evaluations.
EBMs and SBMs each have different strengths. EBMs can calculate Precision, Recall, and F-score, allowing a detailed evaluation of the system performance. In terms of second language acquisition, the evaluation of each edit provides information about the error location, type, and amount, which can improve the quality of feedback and learning efficiency. Most SBMs, on the other hand, can evaluate without references, circumventing the problem of underestimating corrections that are limited by the coverage of references. Also, unlike EBMs, SBMs do not automatically give the lowest score to uncorrected sentences. This allows for a quantifiable measurement to determine whether a sentence has been improved or worsened as a result of correction.
Figure 5: Screenshot of doccano used in the edit-based human evaluation.
@@ 9 Conclusion
To address issues in conventional meta-evaluation in English GEC, we construct a meta-evaluation dataset (SEEDA) consisting of corrections with human ratings along two different evaluation granularities, covering 12 state-of-the-art system corrections including LLMs, and two human corrections with different focuses. The dataset analysis reveals that the results of sentence-level human evaluation differ between granularities and that GEC systems based on GPT and T5 can correct as well as or better than humans. Also, through meta-evaluation using SEEDA, we demonstrate that EBMs may be underestimated in existing meta-evaluations and that matching the evaluation granularity of metrics with human evaluations tends to improve sentence-level correlations. By further analysis, we discovered the uncertainty of conclusions based on a single correlation and found that most metrics lacked the precision to distinguish differences among high-performance neural systems. Finally, we propose a methodology for meta-evaluation and evaluation in GEC. We hope that this paper contributes to further advancements in GEC.
@@ Appendix A Supplement of annotations
Figure 5 shows a screenshot of doccano used in the edit-based human evaluation. The source is enclosed in a <t> tag, and each corrected sentence is emphasized with a <s> tag along with the system number. In step 1, there are error labels for the source and True and False labels for each edit. In step 2, True and False labels with the system number are used to indicate whether the errors in the source were corrected. Due to the specifications of doccano, even if the same edit appears in multiple corrections, annotators need to label each occurrence separately. For information on Appraise in sentence-based human evaluation, you may refer to the Grundkiewicz et al. (2015)’s work.
Acknowledgments
We would like to express our gratitude to the action editor and anonymous reviewers for their constructive feedback. We also thank the annotators who contributed to building the dataset.